OpenStudy (anonymous):

Solving Rational Expressions- Algebra 1 1 3 5 - + - = - 2 t 8

OpenStudy (anonymous):

SO EASY

OpenStudy (anonymous):

T=24

OpenStudy (anonymous):

May I ask how you got that answer?

OpenStudy (anonymous):

ITS T=24

OpenStudy (anonymous):

If your so sure @Stewie_Griffin_ then could you show your work?

OpenStudy (anonymous):

flutter my answer im high, hes right.

OpenStudy (anonymous):

DO U REALLY WANT ME TO EXPLAIN?

OpenStudy (anonymous):

ok, lets get this straight. no worries i will.

OpenStudy (anonymous):

That would be nice, but if its too much to ask then you can leave @Stewie_Griffin_ Gosh.

OpenStudy (anonymous):

U STUPID LMAO

OpenStudy (anonymous):

1/2 + 3/t = 5/8 , 3/t = 5/8 - 1 /2 , 3/t = 1/8 , 1/t = 1/24 , t = 24 .

OpenStudy (anonymous):

CANT EVEN DO SIMPLE MATH

OpenStudy (anonymous):

theres no need to be rude, your obviously in a higher fluttering grade than me @Stewie_Griffin_ and not everyones a fluttering genius BYE

OpenStudy (nurali):

\[\frac{ 1 }{ 2 }+\frac{ 3 }{ t }=\frac{ 5 }{ 8 }\] \[\frac{ 1.t +2.3}{ 2.t }=\frac{ 5 }{ 8 }\] \[\frac{ t+6 }{ 2t }=\frac{ 5 }{ 8 }\] Cross Multiply \[8(t+6)=5(2t)\] \[8t+48=10t\] \[10t-8t=48\] \[2t=48\] \[t=24\]

OpenStudy (anonymous):

THANK YOU @Nurali

OpenStudy (nurali):

Anytime.